OpenStudy (anonymous):

if A = 3 and B = 5 evaluate the expression: 4A - B2

OpenStudy (hoblos):

4A - B2 = 4(3) - 5^2 = 12-25 = -13

OpenStudy (hoblos):

you mean B^2 or B*2 ?

OpenStudy (anonymous):

b^2

OpenStudy (hoblos):

0k thats what i thought so the answer is -13